The probability distribution of X, the number

of imperfections per 10 meters of a synthetic fabric in

continuous rolls of uniform width, is given by

x 01234

f(x) 0.41 0.37 0.16 0.05 0.01

Construct the cumulative distribution function of X.


1
Expert's answer
2022-02-17T04:28:33-0500
"F(0)=P(X\le0)=P(X=0)=0.41"

"F(1)=P(X\le1)=P(X=0\ or\ X=1)"

"=P(X=0)+P(X=1)=0.41+0.37=0.78"

"F(2)=P(X\le2)=P(X=0\ or\ X=1\ or\ X=2)"

"=P(X=0)+P(X=1)+P(X=2)"

"=0.78+0.16=0.94"

"F(3)=P(X\le3)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)=0.94+0.05=0.99"

"F(4)=P(X\le4)=1"

The cumulative distribution of "X" is


"F(x) = \begin{cases}
 0 &x<0 \\
 0.41 & 0\le x<1 \\
 0.78 & 1\le x<2 \\
 0.94 & 2\le x<3 \\
 0.99 & 3\le x<4 \\
 1.00 & x\ge 1
\end{cases}"